Calculate the effective cost of the following loan if the borrower prepays at Loan amount $100,000, Term: 30 years; Interest rate: 75% Monthly Payment, b. 8.645% d. None of the above the end of year 3 8.285% 5%prepayment penalty over c 8.935% 20. You borrow $100,000 choose 30-year term payment between these two mortgages? a. $84,854 b. $102,366 c. $125,786 d. None of the above mortgage with monthly payments. You can either choose 15-year term with interest rate 7% or 1 with Interest rate 8%. If both loans are held to maturity, what is the difference of total interest You borrow S 100,000 at 6% for 30 years with monthly payments. You pay 2 discount points and your APR is 65% What is the amount of your other financing fees besides the discount points? a. $1,144.7 b. $2,144.7 c. $3,144.7 d. None of the above 21 You borrow S 100,000 Constant Arnortization Mortgage (CAM) at 10% for 20 years, assuming annual payment, w your total payment for year 11? a. 5,000 c. 8,000 22 b. 6,000 d. 10,000 (For 23-25) You are buying a house for si 50,000 with a 20% down payment, the lender will finance the remainder onth purchase price with a 30 year CPM with bi-weekly payments at 6.125% annual rate (hint: there are 52 weeks in 23 Approximately how many payments does it take to reduce the loan balance to $100,000? a. 268


May I please have help with 19 and 20 and how to solve them ?
Answer #1

Answer 19:

Correct answer is:

c. 8.935%

Explanation:

Loan amount = $100,000

Term = 30 years = 30 * 12 = 360 months

Interest rate = 7.5%

Monthly interest rate = 7.5%/12

Monthly installment =

PMT (rate, nper, pv, fv, type)

= PMT (7.5%/12, 360, -100000, 0, 0)

= $699.2145

To get outstanding balance at the end of year 3, we will use PV of remaining installments:

Remaining installments = 360 - 3*12 = 324 months

PV (rate, nper, pmt, fv, type)

PV (7.5%/12, 324, -699.2145, 0, 0)

=$97014.25

Hence on prepayment at end of year 3, total amount paid = Balance + Prepayment penalty

= $97014.25 + 5% * $97014.25

= $101,864.96

To calculate effective cost we will use RATE function of excel:

RATE (nper, pmt,pv, fv, type)

= RATE (36, 699.2145, 100000, -101864.96, 0)

= 0.74458%

Monthly cost = 0.74458%

Annual effective cost =0.74458% *12 = 8.935%

Hence option c is correct and other options a, b and d are incorrect.

Answer 20:

Correct answer is:

b. $102,366

Explanation:

Loan amount = 100,000

Mortgage 1:

Monthly Interest rate = 7%/12

Period = 15 years = 15 * 12 = 180

Monthly payment =

PMT (7%/12, 180, -100000, 0,0)

=$898.8283

Total interest payment over the life of loan = Monthly installment * Number of installments - Principal amount

= 898.8283 * 180 - 100000

= $61789.09

Mortgage 2:

Monthly Interest rate = 8%/12

Period = 30 years = 30 * 12 = 360

Monthly payment =

PMT (8%/12, 360, -100000, 0,0)

=$733.7646

Total interest payment over the life of loan = Monthly installment * Number of installments - Principal amount

= 733.7646 * 360 - 100000

= $164,155.26

Hence difference of total interest payment between the two mortgages = 164155.26 - 61789.09 =$102366

Hence option b is correct and other options a, c and d are incorrect.